The height of a soccer ball kicked from the ground follows a path h=- 16t^2+64t where t is the time in seconds. Find a) the maximum height of the soccer ball, b) the time it takes to reach the highest point, and c) the time the ball is in the air.
Answer by richwmiller(17219)

h=- 16t^2+64t
b) vertex
-2/2a=-64/-32=2 seconds to reach max height
a)
h=- 16*2^2+64*2
h=-64+128=64 ft
c)
0=- 16t^2+64t
16t^2=64t
divide by t
16t=64
t=4 seconds in the air